2. If it takes 6 men working 6 hours each to dig a hole in the ground, then how long will it take 1 man to dig a hole that is 1/2 the size of the hole dug by the 6 men in 6 hours?
Answer by venugopalramana(3286)

2. If it takes 6 men working 6 hours each to dig a hole in the ground, then how long will it take 1 man to dig a hole that is 1/2 the size of the hole dug by the 6 men in 6 hours?
MEN................HOLE..............HRS
6....................1................6
1....................1/2..............?
FOR THESE TYPE OF QUESTIONS...
1.PUT THE NUMBER ON THE HEAD OF COLUMN OF QUESTION MARK FIRST FOLLOWED BY MULTIPLICATION SIGN....HERE
6*....
2. NEXT TAKE ONE BY ONE ISSUE..AND THINK HOW IT WILL EFFECT THE ?? OR THE THING TO BE FOUND.IF THE ANSWER SHOULD BE BIGGER PUT HIGHER NUMBER IN THE NUMERATOR AND VICE VERSA.HERE
LET US TKE MEN FIRST....THINK...6 MEN NEED 6 HRS TO DO THE JOB ..SO 1 MAN NEEDS ???MOR HOURS..SO WE HOULD PUT 6 IN N.R...THAT IS 6/1..SO WE GET
6*(6/1)
3...NEXT JOB OF HOLE DIGGING..THINK ..TO DIG 1 OLE WE NEED 6 HRS..TO DIG 1/2 HOLE..WE NEED LESS HRS SO PUT 1/2 IN N.R..SO..
6*(6/1)*{(1/2)/1}=6*6/2=18 HRS IS THE ANSWER